### v1

# ShortGPT: Layers in Large Language Models are More Redundant Than You Expect

Abstract
As Large Language Models (LLMs) continue to advance in performance, their size has escalated significantly, with current LLMs containing billions or even trillions of parameters. However, in this study, we discovered that many layers of LLMs exhibit high similarity, and some layers play a negligible role in network functionality. Based on this observation, we define a metric called Block Influence (BI) to gauge the significance of each layer in LLMs. We then propose a straightforward pruning approach: layer removal, in which we directly delete the redundant layers in LLMs based on their BI scores. Experiments demonstrate that our method, which we call ShortGPT, significantly outperforms previous state-of-the-art (SOTA) methods in model pruning. Moreover, ShortGPT is orthogonal to quantization-like methods, enabling further reduction in parameters and computation. The ability to achieve better results through simple layer removal, as opposed to more complex pruning techniques, suggests a high degree of redundancy in the model architecture.
Figure 1: The results on MMLU by pruning 25% parameters with different methods.
1 Introduction
The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains. Guided by the scaling laws identified in prior work Kaplan et al. (2020); Hoffmann et al. (2022), current LLM research has tended to increase model parameters consistently to boost performance. However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use. To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques Zhu et al. (2023) that reduce the inference cost. Model compression methods can be broadly categorized into two categories: quantization Liu et al. (2021); Gholami et al. (2022); Dettmers et al. (2022, 2024) and pruning LeCun et al. (1989); Han et al. (2015). Quantization methods quantize the weights and activations to lower precision. However, the acceleration benefits of quantization are dependent on hardware support and sometimes require additional fine-tuning to maintain good performance. In contrast, pruning methods remove redundant model parameters to reduce the overall parameter count. Pruning can be directly applied to a trained model without retraining and is generally more hardware-friendly than quantization approaches. While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex. Some require the use of gradient information Zhang et al. (2023) or only focus on compression in width Ashkboos et al. (2024), rendering them overly sophisticated for practical applications. Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models. Identifying effective model pruning techniques necessitates studying model redundancy Huang et al. (2021); Dalvi et al. (2020). Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers. For model pruning in LLMs, most previous works have concentrated on tensor-wise redundancy analysis, investigating the redundancy within each parameter tensor. However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance. For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark Hendrycks et al. (2020) only drop from 55.0 to 52.22 (a 95% retention). Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.19 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model. In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs. We find BI is a more relevant indicator of the layer’s importance in LLMs, and we can perform model pruning by simply deleting redundant layers with BI. This straightforward act of removing specific layers significantly outperforms previous more complex pruning methods. Our findings highlight substantial redundancy in current LLM architectures and shed light on opportunities for more efficient LLM training in the future. The main contributions of our paper are summarized as follows:
  • We analyze the redundancy in large language models (LLMs) and find that they exhibit significant redundancy at the layer level. This insight inspires us to prune LLMs by simply removing redundant layers.
  • We propose a metric called Block Influence (BI) as an effective indicator of layer importance. Through a quantitative analysis, we demonstrate that LLMs possess redundancy both in depth (layers) and width (parameters within layers).
  • Based on the BI metric, we propose a simple yet effective pruning strategy by removing layers with low BI scores. Experimental results show that our method maintains 92% performance while reducing approximately 25% of the parameters and computation, outperforming previous state-of-the-art methods.
  • Furthermore, we demonstrate that our layer pruning approach is orthogonal to quantization methods, meaning it can be combined with quantization techniques to further reduce the deployment overhead of LLMs.
(a) Layer pruning order. We first calculate the BI score of each layer in LLM, and then prune the redundant layer in ascending BI score.
(b) Performance of ShortGPT with increasing pruning ratio.
Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. 2.2), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
2 Methodology
In this section, we present the methodological framework of our layer deletion approach for LLMs, elucidating the underlying principles and techniques employed. We first quantify the layer redundancy problem present in current prominent LLMs, such as LLaMA 2 and Baichuan 2. Then, we introduce a metric called Block Influence (BI), designed to assess the transformation of hidden states by each layer during LLM inference. Building upon BI, we apply layer deletion to LLMs, thereby reducing their inference cost without compromising their predictive accuracy or linguistic capabilities.
2.1 Layer redundancy
The predominant LLMs in current are primarily based on the Transformer Vaswani et al. (2017). The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other. Transformer is a sequence to sequence mapping, which can be defined as $y=f(X,\theta)$, where $X\in R^{T\times n}$, $Y\in R^{T\times n}$, $T$ is the length of the sequence, $n$ is the vocabulary size, $\theta$ is the learnable parameters. The formal expression of an L-layer transformer is as follows:
$\textbf{\footnotesize{X}}_{0}$ $=\textbf{\footnotesize{X}}\textbf{W}_{\text{emb}}$
$\textbf{\footnotesize{A}}_{i+1}$ $=\text{ATTN}(\text{LN}(\textbf{\footnotesize{X}}_{i}))+\textbf{\footnotesize{X}}_{i}$
$\textbf{\footnotesize{X}}_{i+1}$ $=\text{FFN}(\text{LN}(\textbf{\footnotesize{A}}_{i+1}))+\textbf{\footnotesize{A}}_{i+1}$
Y $=\textbf{\footnotesize{X}}_{L}\textbf{W}_{\text{head}}$ (1)
where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ Chowdhery et al. (2023), $d$ is the hidden dim of the transformer. ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers. Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are. Previous works found that the Transformer possesses certain semantic capabilities in earlier layers Hasan et al. (2021). In this work, we uncover a significant level of redundancy between the layers of the transformer.
Layer Redundancy: A network with high redundancy should contain some redundant layers, which have a minimal impact on the network’s final performance. This may be due to these layers having homogenized functionalities compared to other layers in the network. We have found a high degree of redundancy in the current LLMs by omitting specific layers.
(a) Llama2-7B-Base
(b) Baichuan2-7B-Base
Figure 3: Performance of removing single layer.
We observed that in many cases, the network’s prediction of $Y$ remains largely unchanged. This phenomenon demonstrates a significant level of redundancy across the layers of the transformer. Figure 3 shows the perplexity and MMLU Hendrycks et al. (2020) score of omitting a certain layer in the Llama2-7B-Base Touvron et al. (2023), an English based LLMs, and Baichuan2-7B-Base Yang et al. (2023) which is mainly focused on Chinese. For more benchmark results, please refer to the section 3. Figure 3 reveals that some layers do not play a crucial role in LLMs. Dropping a certain layer from the LLMs may have a minimal effect on the final results, and the same holds for the perplexity. Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the final layer often being more critical.
Figure 4: Illustration of our Block Influence(BI) metric. We omit layer norm for visual conciseness.
2.2 Layer importance
As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers. Notably, the deeper layers demonstrate a higher level of redundancy. To remove these redundant layers, we need a model-intrinsic metric to measure the importance of the layer. An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons. In Samragh et al. (2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers. To characterize the influence of a certain layer. We defined a new metric, Block Influence (BI), under the assumption that the more a transformer block changes the hidden states, the more influential this layer is. As Figure 4 shows, the BI score of $i^{th}$ blocks can be calculated as follows:
$\text{BI}_{i}=1-\mathbb{E}_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{||X_{i,t}||_{2}||X_{i+1,t}||_{2}},$ (2)
where $X_{i,t}$ means the $t^{th}$ row of $X_{i}$. Our empirical evidence supports that BI effectively reflects the importance of a layer. Figure 8 shows these different metrics in detail.
2.3 Layer Removal
The method we propose is straightforward: layer removal, which we simply referred as ShortGPT. We delete certain layers in LLMs based on BI score. Figure 2 illustrates our pruning method. First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 Rae et al. (2019). Then we collect the hidden states of each layer during inference on these samples. Next, we calculate the BI score based on the collected hidden states. Finally, we sort layers in ascending order according to the BI, and delete the layers with the smaller importance. The number of layers to be deleted can vary to trade off the speed and performance. The details of our layer remove setting can be found in Appendix B
3 Experiments
3.1 Experimental Setup
3.1.1 Models
To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B Touvron et al. (2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B. They are all large language models based on the decoder-only Transformer architecture. LLaMA 2 was trained on more than 2 trillion tokens. Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE Su et al. (2024) positional embedding with ALiBi Press et al. (2021).
3.1.2 Benchmarks
In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU Hendrycks et al. (2020), CMMLU Li et al. (2024) for evaluating large models. In addition, we also followed LaCo Yang et al. (2024) to evaluate a wider dataset.
MMLU Hendrycks et al. (2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings. This makes benchmarks more challenging and similar to the way we evaluate humans. This benchmark covers 57 subjects including STEM, humanities, social sciences, etc. Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
CMMLU Li et al. (2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture. CMMLU covers 67 topics, from elementary school to university or professional level. Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
CMNLI Xu et al. (2020) is part of the Chinese language understanding assessment benchmark. It consists of two parts: XNLI and MNLI. HellaSwag (HeSw) Zellers et al. (2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans. PIQA Bisk et al. (2020) is a multi-choice question and answer dataset that focuses on daily scenarios. This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios. CHID Zheng et al. (2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms. CoQA Reddy et al. (2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers. BoolQ Clark et al. (2019) is a question-answer dataset containing 15942 examples of yes/no questions. These problems occur naturally - they are generated in an environment that is silent and unconstrained. Race Lai et al. (2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students. XSumHasan et al. (2021) is used to evaluate abstract single document summarization systems. The goal is to create a short, one-sentence new summary of what the article is about. C3 Sun et al. (2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams. PG19 Rae et al. (2019) is a long document dataset from books used to test the effectiveness of language modeling.
3.1.3 Baselines
To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo Yang et al. (2024). For our method, we use PG19 for layer importance and perplexity calculation.
LLMPru. Ma et al. (2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
SliceGPT Ashkboos et al. (2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
LaCo Yang et al. (2024) is a pruning method for large language models based on merging multiple layers.
LLM Method Ratio Benchmarks Ave. Per.
CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU
Llama2-7B Dense 0.00% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92 44.52 100.00
LLMPrun. 27.0% 34.33 56.46 71.22 25.25 36.54 42.51 55.20 22.56 22.35 11.51 25.64 23.33 25.25 32.84 73.76
SliceGPT 26.4% 31.70 50.27 66.21 20.79 36.54 41.36 38.32 21.07 21.66 4.89 39.78 28.92 25.37 32.84 73.76
LaCo 27.1% 34.43 55.69 69.80 36.14 40.38 45.70 64.07 22.61 23.61 15.64 39.67 26.45 25.24 38.41 86.28
ShortGPT 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.71 32.25 35.17 0.67 39.62 43.96 32.25 42.60 95.69
Llama2-13B Dense 0.00% 32.99 74.78 79.71 47.35 50.00 66.91 82.39 57.95 60.38 23.45 47.51 55.00 38.40 51.91 100.00
LLMPrun. 24.4% 33.03 67.76 76.66 35.64 40.38 50.86 56.42 22.47 22.08 19.17 32.33 25.21 24.71 38.97 75.07
SliceGPT 23.6% 29.82 55.71 69.04 19.31 36.54 47.26 37.86 23.41 24.03 5.27 41.92 37.14 25.79 34.84 67.11
LaCo 24.6% 32.86 64.39 74.27 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 48.30 93.05
ShortGPT 24.6% 32.99 56.14 68.93 22.93 54.61 52.42 71.80 53.92 57.45 2.14 39.01 52.42 35.36 48.10 92.95
Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 26.96 24.09 20.82 64.55 53.87 56.95 53.63 100.00
LLMPrun. 24.2% 32.28 53.66 71.82 69.80 53.85 47.83 61.19 21.96 22.28 15.98 41.64 24.93 25.69 41.76 77.87
SliceGPT 22.2% 32.07 25.29 50.33 14.85 36.54 19.57 39.30 23.53 22.49 0.00 26.58 25.18 25.25 26.23 56.38
LaCo 24.2% 33.00 52.28 68.50 76.24 42.31 47.26 56.15 28.99 27.72 12.03 50.85 31.53 31.24 42.93 80.05
ShortGPT 24.2% 33.30 56.96 67.68 65.63 50.00 46.70 67.83 53.26 46.76 0.04 56.33 45.77 47.87 49.08 91.52
Baichuan2-13B Dense 0.00% 33.21 71.10 78.07 86.51 50.00 65.6 77.89 67.27 68.94 25.02 65.64 59.50 61.30 62.31 100.00
LLMPrun. 24.3% 33.80 53.57 71.82 72.77 37.50 38.82 56.54 21.17 21.61 13.67 39.89 23.19 25.18 39.20 62.91
SliceGPT 22.8% 32.07 25.85 51.03 10.40 36.54 18.02 37.83 21.56 21.52 0.00 24.99 22.95 25.26 25.03 40.17
LaCo 24.7% 33.03 60.71 68.88 76.73 44.23 55.45 62.35 56.92 57.80 12.32 61.10 51.35 53.65 53.43 85.75
ShortGPT 24.7% 32.81 60.55 71.60 80.17 47.13 54.30 62.54 55.77 56.41 15.14 60.16 52.11 58.86 54.43 87.33
Table 1: Comparison of pruning methods on multiple natural language benchmarks. The results of LLMPrun., SliceGPT and LaCo are reported from LaCo.
3.2 Main Results
To validate the efficacy of our proposed method, we conducted comparative experiments against benchmark and baseline techniques commonly employed in large language model evaluation. Considering the current structured pruning methods generally reduce parameters by no more than 30%, we performed experiments with approximately 1/4 of the parameters pruned. The experimental results are presented in Table 1. Additional experiments exploring different parameter reduction proportions will be discussed in the subsequent section.
The results demonstrate that the performance of the model pruned by our method significantly surpasses that of the baseline method, maintaining most of the large language model’s capabilities, such as reasoning, language understanding, knowledge retention, and examination performance, to the greatest extent possible. Furthermore, we note that the approach of reducing the number of layers (ShortGPT/LaCo) outperforms the method of reducing the embedding dimensions (LLMPru./SliceGPT), implying that the model exhibits more redundancy in-depth than in width. Further experimental analysis will be presented in the ensuing section.
4 Analysis
4.1 Varying prune ratio
Figure 5: Pruning by different metrics, with increasing pruning ratio.
As a pruning method, we further validated the effects of different pruning ratios on model performance. Experiments were conducted on the Llama2-7B-Base and Baichuan2-7B-Base models, observing the Perplexity and MMLU. The pruning ratios ranged from 0% to 97%, employing the strategy of deleting layers based on the importance proposed in this paper, resulting in models with 1 to 32 layers. The results in Figure 5 indicate that as the pruning ratio increases, the performance of the model declines. However, the MMLU score drops significantly at a certain layer, which may imply the existence of certain special layers within the network that play a critically important role.
4.2 Depth Redundancy vs width Redundancy
Figure 6: Redundancy on width, removing a single head from model. The x-axis is the head id.
In the preceding sections, we analyzed and explored the redundancy of large language models (LLMs) in terms of depth (layers). However, we also investigated redundancy in terms of width, specifically focusing on the attention heads. Our method involves sequentially removing each head from the multi-head attention mechanism and adjusting the parameters of the attention output projection to ensure the output shape of the Attention block remains unchanged. The Feed-Forward Network (FFN) remains unaltered. Similar to the approach in Section 2.1, we observe the impact of each head’s removal on the final performance of the network. Figure 6 illustrates the changes in Perplexity and MMLU scores after each head is removed. The results indicate that LLMs exhibit a high degree of redundancy in width, comparable to their redundancy in depth. However, this width redundancy does not follow any discernible pattern and varies across different models. We hypothesize that this variation may be attributed to the symmetry among heads.
4.3 Redundancy on non-transformer LLM
Figure 7: Redundancy on RWKV, removing single layers on RWKV-7B.
We further investigate whether the redundancy observed in depth is a result of the Transformer’s architecture. Given that the vast majority of popular LLMs are based on the Transformer structure. We selected the RWKV-7B model proposed in Peng et al. (2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent. We analyzed this model using the methodology from 2.1. Figure 7 shows the redundancy of the RWKV-7B 111 We use rwkv-v5-world-7B from https://huggingface.co/BlinkDL/rwkv-5-world model. Through this figure, we can observe that the RWKV-7B model also exhibits a high level of redundancy. This may suggest that the redundancy is universal across LLMs.
4.4 Importance metric
Figure 8: Different importance metrics.
The overarching concept of our method is to rank layers by importance and remove the less significant ones. The definition of importance metrics has a profound impact on the outcome. In this section, we define and compare various important metrics:
Sequential: The importance is directly proportional to the sequence order, with shallower layers being more important. This can be implemented by taking the negative of each layer’s index as the importance metric.
Reverse-order/Norm: This metric posits that importance is inversely proportional to the sequence order. It assigns higher importance scores to the shallower layers Inspired by Figure 3. This method gives the same order as measuring importance by hidden states norm mentioned in 2.2. Here, the importance metric is the layer’s index itself.
Relative Magnitude: Proposed in Samragh et al. (2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
BI: BI mentioned in previous section 4.4.
Figure 8 demonstrates the different metrics. We observe that shallower layers in the LLM network are more crucial than deeper ones. Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics. The method of Samragh et al. (2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent. It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation. This also indicates the limitations of the current LLMs evaluation methods.
4.5 Orthogonal to Quantization
In this section, we show that our method is orthogonal to quantization methods. We apply our method to a model  [^2] We take the model from https://huggingface.co/TheBloke/Llama-2-7B-GPTQ quantized by GPTQ algorithm. Table 2 shows that our method is compatible with the quantization-like method, which could reduce the footprint further.
Model Ratio/Layer Perplexity MMLU
Baseline 0%/32 8.4999 37.99
ShortGPT 27.1%/23 42.6951 36.69
Table 2: Layer removal results on Llama2-7B-Base-GPTQ.
5 Limitation
Although our method demonstrates strong competitiveness compared to current pruning methods, with some benchmarks even suggesting that deleting certain layers has no impact on the final result, this does not imply that layer removal comes without drawbacks. Due to the limitations of benchmarking, current evaluations may not fully capture the impact of layer removal on model performance. Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks. On benchmarks such as GSM8K Cobbe et al. (2021) and HumanEval Chen et al. (2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero. Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal. Post-training techniques could potentially further mitigate performance losses, which is an area of ongoing research.
6 Related works
To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into four categories: model pruning, knowledge distillation, quantization, and low-rank factorization.
Model pruning: model pruning LeCun et al. (1989); Han et al. (2015) is a classic and effective method of reducing model redundancy modules to compress models. The model pruning methods mainly include unstructured pruning and structured pruning. The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT Frantar and Alistarh (2023) and LoRAPrune Zhang et al. (2023). However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition. Another more practical approach is structured pruning, GUMSyed et al. (2023) makes an analysis of several structured pruning methods for decoder-only LLMs. LLM-Pruner Ma et al. (2024) selectively removes non-critical structures according to gradient information. ShearedLLaMA Xia et al. (2023) employs targeted structured pruning and dynamic batch loading. LaCo Yang et al. (2024) used layer merging to compress the model. Compared to the previous method, our method is a simple and efficient structured pruning method.
Knowledge distillation: Knowledge distillation (KD) Hinton et al. (2015); Gou et al. (2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network. Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results Li et al. (2022). In addition, when we can access transparent teacher models, the result can be improved more. MiniLLM Gu et al. (2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution. DistlLLM Ko et al. (2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs. However, compared to model pruning, this method often requires higher computational resources.
Quantization: quantization Liu et al. (2021); Gholami et al. (2022); Dettmers et al. (2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models. Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms. LUT-GEMM Park et al. (2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format. SPQR Dettmers et al. (2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits. Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
Low-rank factorization: low-rank decomposition Cheng et al. (2017); Povey et al. (2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions. TensorGPT Xu et al. (2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices. Recently, SliceGPT Ashkboos et al. (2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
Model redundancy: researchers have long noticed the significant redundancy in nonlinear models Catchpole and Morgan (1997). In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy. In Bian et al. (2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads. In Dalvi et al. (2020), researchers dissect two pre-trained models, BERT Devlin et al. (2018) and XLNet Yang et al. (2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level. However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
7 Conclusion
This work has introduced a novel approach to pruning Large Language Models (LLMs) based on layer redundancy and an "importance" metric defined as attention entropy. Our study reveals a significant degree of layer-wise redundancy in LLMs, indicating that certain layers contribute minimally to the overall network function and can thus be removed without substantially compromising model performance. By employing a straightforward layer removal strategy guided by the calculated importance of each layer, we have demonstrated that it is possible to maintain up to 95% of an LLM’s performance while reducing the model’s parameter count and computational requirements by approximately 25%. This achievement not only surpasses previous state-of-the-art pruning methods but also underscores the potential for further optimizations in LLM deployment strategies.
Our findings suggest that the redundancy inherent in LLMs is largely depth-based rather than width-based, highlighting an avenue for future research into the structural efficiency of neural networks. Moreover, our pruning approach, characterized by its simplicity and effectiveness, is compatible with other compression techniques such as quantization, offering a composite pathway to model size reduction that is both significant and versatile.
The implications of our research extend beyond academic interest, offering practical solutions for deploying advanced LLMs in resource-constrained environments. By enabling more efficient model architectures without the need for extensive retraining, our pruning method facilitates broader access to cutting-edge AI capabilities across a variety of platforms and devices.
In conclusion, our investigation into layer redundancy and the development of an importance-based pruning strategy represent a meaningful advancement in the optimization of Large Language Models. As the demand for sophisticated AI tools continues to grow, approaches like ours will play a crucial role in making these technologies more accessible and sustainable. Future work will focus on refining our understanding of model redundancy and exploring additional methods for enhancing the efficiency of neural network models.
Appendix A Detailed Strategy for Layer Removal
Table 3: Strategies for Layer Removal in Models.
Strategy
Description
Sequential
Layers are removed sequentially from the beginning of the model. The process starts with layer 0 and progressively includes more layers for removal (e.g., {0}, {0, 1}, …).
Reverse-order
This strategy involves starting from the model’s final layer and progressively removing layers in reverse order (e.g., {-1}, {-1, -2}, …).
Relative Magnitude
Layers are removed in ascending order based on their Relative Magnitude values. The removal process accumulates layers from those with the smallest to the largest values, mirroring the sequential strategy’s accumulation method.
BI (Block Influence)
Follows a similar accumulation approach as the Sequential strategy, but layers are ordered and removed according to their BI values, starting from the lowest and moving to the highest.
Appendix B Benchmark Setup for Removed Layers
Before delineating the specific layer removal configurations in the benchmarked models, a concise overview of the structural modifications is warranted. The Llama-2-7B and Baichuan-2-7B models, initially designed with 32 layers, have been streamlined by excising 9 layers, corresponding to a 28% reduction of the total layer framework. The 13B variants, originally constituted of 40 layers, experienced a decrement of 10 layers, marking a 25% diminution. The selection criterion for layer removal was based on ascending Block Influence (BI) values, ensuring a data-driven approach to minimization. This premise underpins the benchmark configurations.
Table 4: Setup of Removed Layers for Benchmark Models
Model Removed Layers
Llama-2-7B 27, 26, 25, 28, 24, 29, 23, 21, 22
Llama-2-13B 33, 31, 32, 30, 29, 34, 28, 35, 27, 26
Baichuan-2-7B 26, 27, 25, 28, 24, 29, 23, 22, 30
Baichuan-2-13B 32, 31, 33, 30, 34, 29, 28, 35, 27, 26
Appendix C Layer Removal on Baichuan2-series Model
Figure 9: Pruning by different metrics on Baichuan2-series model.
Appendix D Layer Removal on HumanEval.
We also found that layer removal has a significant impact on the performance of the model in generative tasks. Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval Chen et al. (2021), which is a dataset used to evaluate the performance of code generation models. The results are shown in Figure 10. After removing 10 layers, the performance on HumanEval almost completely degraded. We speculate that it may be due to significant cumulative errors in generative tasks.
Figure 10: Performance degragation on generation task.

### v2

@@
 The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains.
-Guided by the scaling laws identified in prior work Kaplan et al.
-(2020); Hoffmann et al.
-(2022), current LLM research has tended to increase model parameters consistently to boost performance.
+Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research has tended to increase model parameters consistently to boost performance.
 However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use.
-To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques Zhu et al.
-(2023) that reduce the inference cost.
-Model compression methods can be broadly categorized into two categories: quantization Liu et al.
-(2021); Gholami et al.
-(2022); Dettmers et al.
-(2022, 2024) and pruning LeCun et al.
-(1989); Han et al.
-(2015).
+To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques (Zhu et al., 2023) that reduce the inference cost.
+Model compression methods can be broadly categorized into two categories: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) and pruning (LeCun et al., 1989; Han et al., 2015).
 Quantization methods quantize the weights and activations to lower precision.
@@
 While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex.
-Some require the use of gradient information Zhang et al.
-(2023) or only focus on compression in width Ashkboos et al.
-(2024), rendering them overly sophisticated for practical applications.
+Some require the use of gradient information (Zhang et al., 2023) or only focus on compression in width (Ashkboos et al., 2024), rendering them overly sophisticated for practical applications.
 Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models.
-Identifying effective model pruning techniques necessitates studying model redundancy Huang et al.
-(2021); Dalvi et al.
-(2020).
+Identifying effective model pruning techniques necessitates studying model redundancy (Huang et al., 2021; Dalvi et al., 2020).
 Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers.
@@
 However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance.
-For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark Hendrycks et al.
-(2020) only drop from 55.0 to 52.22 (a 95% retention).
-Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.19 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model.
+For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark (Hendrycks et al., 2020) only drop from 55.0 to 52.2 (a 95% retention).
+Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.2 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model.
 In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs.
@@
 (b) Performance of ShortGPT with increasing pruning ratio.
-Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. 2.2), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
+Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. LABEL:layerimportacne), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
 2 Methodology
@@
 2.1 Layer redundancy
-The predominant LLMs in current are primarily based on the Transformer Vaswani et al.
-(2017).
+The predominant LLMs in current are primarily based on the Transformer (Vaswani et al., 2017).
 The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other.
@@
 Y $=\textbf{\footnotesize{X}}_{L}\textbf{W}_{\text{head}}$ (1)
-where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ Chowdhery et al.
-(2023), $d$ is the hidden dim of the transformer.
+where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ (Chowdhery et al., 2023), $d$ is the hidden dim of the transformer.
 ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers.
 Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are.
-Previous works found that the Transformer possesses certain semantic capabilities in earlier layers Hasan et al.
-(2021).
+Previous works found that the Transformer possesses certain semantic capabilities in earlier layers (Hasan et al., 2021).
 In this work, we uncover a significant level of redundancy between the layers of the transformer.
@@
 This phenomenon demonstrates a significant level of redundancy across the layers of the transformer.
-Figure 3 shows the perplexity and MMLU Hendrycks et al.
-(2020) score of omitting a certain layer in the Llama2-7B-Base Touvron et al.
-(2023), an English based LLMs, and Baichuan2-7B-Base Yang et al.
-(2023) which is mainly focused on Chinese.
+Figure 3 shows the perplexity and MMLU (Hendrycks et al., 2020) score of omitting a certain layer in the Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese.
 For more benchmark results, please refer to the section 3.
@@
 An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons.
-In Samragh et al.
-(2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers.
+In (Samragh et al., 2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers.
 To characterize the influence of a certain layer.
@@
 Figure 2 illustrates our pruning method.
-First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 Rae et al.
-(2019).
+First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 (Rae et al., 2019).
 Then we collect the hidden states of each layer during inference on these samples.
@@
 3.1.1 Models
-To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B Touvron et al.
-(2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B.
+To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B (Touvron et al., 2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B.
 They are all large language models based on the decoder-only Transformer architecture.
 LLaMA 2 was trained on more than 2 trillion tokens.
-Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE Su et al.
-(2024) positional embedding with ALiBi Press et al.
-(2021).
+Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE (Su et al., 2024) positional embedding with ALiBi (Press et al., 2021).
 3.1.2 Benchmarks
-In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU Hendrycks et al.
-(2020), CMMLU Li et al.
-(2024) for evaluating large models.
-In addition, we also followed LaCo Yang et al.
-(2024) to evaluate a wider dataset.
-MMLU Hendrycks et al.
-(2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings.
+In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024) for evaluating large models.
+In addition, we also followed LaCo (Yang et al., 2024) to evaluate a wider dataset.
+MMLU (Hendrycks et al., 2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings.
 This makes benchmarks more challenging and similar to the way we evaluate humans.
@@
 Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
-CMMLU Li et al.
-(2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture.
+CMMLU (Li et al., 2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture.
 CMMLU covers 67 topics, from elementary school to university or professional level.
 Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
-CMNLI Xu et al.
-(2020) is part of the Chinese language understanding assessment benchmark.
+CMNLI (Xu et al., 2020) is part of the Chinese language understanding assessment benchmark.
 It consists of two parts: XNLI and MNLI.
-HellaSwag (HeSw) Zellers et al.
-(2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans.
-PIQA Bisk et al.
-(2020) is a multi-choice question and answer dataset that focuses on daily scenarios.
+HellaSwag (HeSw) (Zellers et al., 2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans.
+PIQA (Bisk et al., 2020) is a multi-choice question and answer dataset that focuses on daily scenarios.
 This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios.
-CHID Zheng et al.
-(2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms.
-CoQA Reddy et al.
-(2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers.
-BoolQ Clark et al.
-(2019) is a question-answer dataset containing 15942 examples of yes/no questions.
+CHID (Zheng et al., 2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms.
+CoQA (Reddy et al., 2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers.
+BoolQ (Clark et al., 2019) is a question-answer dataset containing 15942 examples of yes/no questions.
 These problems occur naturally - they are generated in an environment that is silent and unconstrained.
-Race Lai et al.
-(2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students.
-XSumHasan et al.
-(2021) is used to evaluate abstract single document summarization systems.
+Race (Lai et al., 2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students.
+XSum(Hasan et al., 2021) is used to evaluate abstract single document summarization systems.
 The goal is to create a short, one-sentence new summary of what the article is about.
-C3 Sun et al.
-(2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams.
-PG19 Rae et al.
-(2019) is a long document dataset from books used to test the effectiveness of language modeling.
+C3 (Sun et al., 2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams.
+PG19 (Rae et al., 2019) is a long document dataset from books used to test the effectiveness of language modeling.
 3.1.3 Baselines
-To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo Yang et al.
-(2024).
+To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo (Yang et al., 2024).
 For our method, we use PG19 for layer importance and perplexity calculation.
-LLMPru. Ma et al. (2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
-SliceGPT Ashkboos et al. (2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
-LaCo Yang et al. (2024) is a pruning method for large language models based on merging multiple layers.
+LLMPru.
+(Ma et al., 2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
+LLMPru.
+applies post training to the pruned model, but for fair comparison, we do not apply post training to it.
+SliceGPT (Ashkboos et al., 2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
+Specifically, they applied PCA to the hidden representation from shallow to deep layers, and incorporated the dimension reduction matrix into existing network parameters.
+LaCo (Yang et al., 2024) is a pruning method for large language models based on reducing layers.
+LaCo gradually merges similar layers from deep to shallow and sets a threshold to avoid continuously merging too many layers.
 LLM Method Ratio Benchmarks Ave. Per.
@@
 LaCo 24.6% 32.86 64.39 74.27 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 48.30 93.05
-ShortGPT 24.6% 32.99 56.14 68.93 22.93 54.61 52.42 71.80 53.92 57.45 2.14 39.01 52.42 35.36 48.10 92.95
+ShortGPT 24.6% 33.00 66.64 73.45 36.61 50.00 58.64 62.48 58.35 60.17 17.59 46.90 54.69 38.38 50.53 97.34
 Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 26.96 24.09 20.82 64.55 53.87 56.95 53.63 100.00
@@
 However, the MMLU score drops significantly at a certain layer, which may imply the existence of certain special layers within the network that play a critically important role.
-4.2 Depth Redundancy vs width Redundancy
+4.2 Depth Redundancy v.s. Width Redundancy
 Figure 6: Redundancy on width, removing a single head from model. The x-axis is the head id.
@@
 Given that the vast majority of popular LLMs are based on the Transformer structure.
-We selected the RWKV-7B model proposed in Peng et al.
-(2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent.
+We selected the RWKV-7B model proposed in (Peng et al., 2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent.
 We analyzed this model using the methodology from 2.1.
@@
 4.4 Importance metric
-Figure 8: Different importance metrics.
+Figure 8: Different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer
 The overarching concept of our method is to rank layers by importance and remove the less significant ones.
@@
 It assigns higher importance scores to the shallower layers Inspired by Figure 3.
-This method gives the same order as measuring importance by hidden states norm mentioned in 2.2.
+This method gives the same order as measuring importance by hidden states norm mentioned in LABEL:layerimportacne.
 Here, the importance metric is the layer’s index itself.
-Relative Magnitude: Proposed in Samragh et al. (2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
+Relative Magnitude: Proposed in (Samragh et al., 2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
 BI: BI mentioned in previous section 2.2.
@@
 Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics.
-The method of Samragh et al.
-(2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent.
+The method of (Samragh et al., 2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent.
 It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation.
@@
 Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks.
-On benchmarks such as GSM8K Cobbe et al.
-(2021) and HumanEval Chen et al.
-(2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero.
+On benchmarks such as GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero.
 Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal.
@@
 To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into four categories: model pruning, knowledge distillation, quantization, and low-rank factorization.
-Model pruning: model pruning LeCun et al.
-(1989); Han et al.
-(2015) is a classic and effective method of reducing model redundancy modules to compress models.
+Model pruning: model pruning (LeCun et al., 1989; Han et al., 2015) is a classic and effective method of reducing model redundancy modules to compress models.
 The model pruning methods mainly include unstructured pruning and structured pruning.
-The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT Frantar and Alistarh (2023) and LoRAPrune Zhang et al.
-(2023).
+The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar and Alistarh, 2023) and LoRAPrune (Zhang et al., 2023).
 However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition.
-Another more practical approach is structured pruning, GUMSyed et al.
-(2023) makes an analysis of several structured pruning methods for decoder-only LLMs.
-LLM-Pruner Ma et al.
-(2024) selectively removes non-critical structures according to gradient information.
-ShearedLLaMA Xia et al.
-(2023) employs targeted structured pruning and dynamic batch loading.
-LaCo Yang et al.
-(2024) used layer merging to compress the model.
+Another more practical approach is structured pruning, GUM(Syed et al., 2023) makes an analysis of several structured pruning methods for decoder-only LLMs.
+LLM-Pruner (Ma et al., 2024) selectively removes non-critical structures according to gradient information.
+ShearedLLaMA (Xia et al., 2023) employs targeted structured pruning and dynamic batch loading.
+LaCo (Yang et al., 2024) used layer merging to compress the model.
 Compared to the previous method, our method is a simple and efficient structured pruning method.
-Knowledge distillation: Knowledge distillation (KD) Hinton et al.
-(2015); Gou et al.
-(2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network.
-Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results Li et al.
-(2022).
+Knowledge distillation: Knowledge distillation (KD) (Hinton et al., 2015; Gou et al., 2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network.
+Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results (Li et al., 2022).
 In addition, when we can access transparent teacher models, the result can be improved more.
-MiniLLM Gu et al.
-(2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution.
-DistlLLM Ko et al.
-(2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs.
+MiniLLM (Gu et al., 2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution.
+DistlLLM (Ko et al., 2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs.
 However, compared to model pruning, this method often requires higher computational resources.
-Quantization: quantization Liu et al.
-(2021); Gholami et al.
-(2022); Dettmers et al.
-(2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models.
+Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models.
 Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms.
-LUT-GEMM Park et al.
-(2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format.
-SPQR Dettmers et al.
-(2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits.
+LUT-GEMM (Park et al., 2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format.
+SPQR (Dettmers et al., 2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits.
 Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
-Low-rank factorization: low-rank decomposition Cheng et al.
-(2017); Povey et al.
-(2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions.
-TensorGPT Xu et al.
-(2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices.
-Recently, SliceGPT Ashkboos et al.
-(2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
-Model redundancy: researchers have long noticed the significant redundancy in nonlinear models Catchpole and Morgan (1997).
+Low-rank factorization: low-rank decomposition (Cheng et al., 2017; Povey et al., 2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions.
+TensorGPT (Xu et al., 2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices.
+Recently, SliceGPT (Ashkboos et al., 2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
+Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole and Morgan, 1997).
 In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy.
-In Bian et al.
-(2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads.
-In Dalvi et al.
-(2020), researchers dissect two pre-trained models, BERT Devlin et al.
-(2018) and XLNet Yang et al.
-(2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level.
+In (Bian et al., 2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads.
+In (Dalvi et al., 2020), researchers dissect two pre-trained models, BERT (Devlin et al., 2018) and XLNet (Yang et al., 2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level.
 However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
@@
 We also found that layer removal has a significant impact on the performance of the model in generative tasks.
-Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval Chen et al.
-(2021), which is a dataset used to evaluate the performance of code generation models.
+Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval (Chen et al., 2021), which is a dataset used to evaluate the performance of code generation models.
 The results are shown in Figure 10.

### v3

@@
 Abstract
-As Large Language Models (LLMs) continue to advance in performance, their size has escalated significantly, with current LLMs containing billions or even trillions of parameters.
-However, in this study, we discovered that many layers of LLMs exhibit high similarity, and some layers play a negligible role in network functionality.
-Based on this observation, we define a metric called Block Influence (BI) to gauge the significance of each layer in LLMs.
-We then propose a straightforward pruning approach: layer removal, in which we directly delete the redundant layers in LLMs based on their BI scores.
-Experiments demonstrate that our method, which we call ShortGPT, significantly outperforms previous state-of-the-art (SOTA) methods in model pruning.
+As Large Language Models (LLMs) continue to advance in performance, their size has increased significantly, with current LLMs containing billions or even trillions of parameters.
+In this study, we identify notable redundancy across the layers of LLMs, where some layers contribute minimally to overall network functionality.
+To quantify this, we introduce a metric called Block Influence (BI) which use the similarity between layer’s input and output to measure the importance of each layer.
+Based on the observation of layer redundancy, we propose a straightforward pruning method: layer removal, which eliminates redundant layers based on their BI scores.
+Our approach, termed ShortGPT, demonstrates superior performance over previous state-of-the-art pruning methods.
 Moreover, ShortGPT is orthogonal to quantization-like methods, enabling further reduction in parameters and computation.
-The ability to achieve better results through simple layer removal, as opposed to more complex pruning techniques, suggests a high degree of redundancy in the model architecture.
-Figure 1: The results on MMLU by pruning 25% parameters with different methods.
+The ability to achieve better results through simple layer removal, as opposed to more complex pruning techniques, suggests a high degree of redundancy across layers, not only in transformer models but also in non-transformer models.
+We hope this work will contribute to future research in LLM compression.
 1 Introduction
 The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains.
-Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research has tended to increase model parameters consistently to boost performance.
-However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use.
-To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques (Zhu et al., 2023) that reduce the inference cost.
-Model compression methods can be broadly categorized into two categories: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) and pruning (LeCun et al., 1989; Han et al., 2015).
-Quantization methods quantize the weights and activations to lower precision.
-However, the acceleration benefits of quantization are dependent on hardware support and sometimes require additional fine-tuning to maintain good performance.
-In contrast, pruning methods remove redundant model parameters to reduce the overall parameter count.
-Pruning can be directly applied to a trained model without retraining and is generally more hardware-friendly than quantization approaches.
-While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex.
-Some require the use of gradient information (Zhang et al., 2023) or only focus on compression in width (Ashkboos et al., 2024), rendering them overly sophisticated for practical applications.
-Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models.
-Identifying effective model pruning techniques necessitates studying model redundancy (Huang et al., 2021; Dalvi et al., 2020).
-Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers.
-For model pruning in LLMs, most previous works have concentrated on tensor-wise redundancy analysis, investigating the redundancy within each parameter tensor.
-However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance.
-For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark (Hendrycks et al., 2020) only drop from 55.0 to 52.2 (a 95% retention).
-Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.2 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model.
-In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs.
-We find BI is a more relevant indicator of the layer’s importance in LLMs, and we can perform model pruning by simply deleting redundant layers with BI.
-This straightforward act of removing specific layers significantly outperforms previous more complex pruning methods.
-Our findings highlight substantial redundancy in current LLM architectures and shed light on opportunities for more efficient LLM training in the future.
+Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research tend to increase model parameters to boost performance.
+As a result, modern LLMs, which can comprise billions to trillions of parameters, require significant hardware resources for deployment, creating substantial barriers to their practical use.
+To mitigate the hardware demands of large models, model compression techniques have become a critical area of focus (Zhu et al., 2023).
+These techniques are generally divided into quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022; 2024) and pruning(LeCun et al., 1989; Han et al., 2015; Frantar & Alistarh, 2023).
+Quantization reduces the precision of model parameters, but its effectiveness often requires specific hardware support.
+In contrast, pruning method removes redundant parameters to decrease the model’s size and computation, offering a more flexible and hardware-agnostic approach.
+Despite its advantages, many existing pruning methods are complex; for example, some require gradient information (Ma et al., 2024), which limits their practicality.
+In this paper, we focus on the issue of layer redundancy in LLMs and propose a novel approach for simplifying these models.
+We introduce Block Influence (BI), a metric that quantifies how much the hidden state changes after passing through each layer, providing a more direct measure of a layer’s importance.
+Leveraging this insight, we propose a simple yet effective pruning method ShortGPT, which identifies and removes layers with lower BI scores, significantly reducing model size without sacrificing much performance.
+To evaluate our approach, we conducted evaluation across comprehensive benchmarks.
+Our experiments revealed that our method exhibits a smaller performance decrement compared to the previous methods.
+For instance, removing 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model resulted in only a slight drop in performance on the MMLU benchmark (Hendrycks et al., 2020), from 55.0 to 52.2.
+Our findings highlight substantial redundancy in current LLMs and suggest potential avenues for improving the efficiency of model training by reducing inherent redundancy in the future.
 The main contributions of our paper are summarized as follows:
   • We analyze the redundancy in large language models (LLMs) and find that they exhibit significant redundancy at the layer level.
-This insight inspires us to prune LLMs by simply removing redundant layers.
-  • We propose a metric called Block Influence (BI) as an effective indicator of layer importance.
-Through a quantitative analysis, we demonstrate that LLMs possess redundancy both in depth (layers) and width (parameters within layers).
-  • Based on the BI metric, we propose a simple yet effective pruning strategy by removing layers with low BI scores.
-Experimental results show that our method maintains 92% performance while reducing approximately 25% of the parameters and computation, outperforming previous state-of-the-art methods.
+This finding inspire us to prune LLMs by simply removing redundant layers.
+  • We propose a metric called Block Influence (BI) as an indicator of layer importance.
+Based on BI, our layer removal method maintains approximately 90% performance while reducing approximately 25% of parameters, outperforming previous state-of-the-art methods.
   • Furthermore, we demonstrate that our layer pruning approach is orthogonal to quantization methods, meaning it can be combined with quantization techniques to further reduce the deployment overhead of LLMs.
-(a) Layer pruning order. We first calculate the BI score of each layer in LLM, and then prune the redundant layer in ascending BI score.
-(b) Performance of ShortGPT with increasing pruning ratio.
-Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. LABEL:layerimportacne), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
+(a) Perplxity
+(b) MMLU
+Figure 1: Performance of removing certain layer from LLMs. We can see that certain layers are redundant, and their removal results in minimal performance degradation.
+2 Motivation
+2.1 Background
+The predominant LLMs are primarily based on the Transformer architecture (Vaswani et al., 2017), with the pre-norm configuration being the most commonly adopted, as in models like LLaMA (Touvron et al., 2023).
+The pre-norm configuration, where layer normalization is applied before the self-attention and feed-forward layers, offers several advantages such as faster convergence, improved training stability, and better scalability for deeper networks (Xiong et al., 2020; Liu et al., 2020; Wang et al., 2024).
+Due to these benefits, the pre-norm approach has been adopted even in non-transformer models, such as Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023).
+For the sake of simplicity in descriptions, our analysis primarily focuses on the Transformer architecture, though we extend our experiments to non-Transformer structures in Section 4.4.
+Figure 2: The cosine similarity between a layer’s input and output during the training process.
+The horizontal axis (X-axis) represents the number of training tokens, while the vertical axis (Y-axis) depicts the degree of similarity.
+Notably, the model employing post-normalization exhibits divergence after approximately $\sim$26B tokens of training.
+Training setting is provided in E.
+However, we observe that when pre-norm is adopted, the similarity between the input and output of transformer layers tends to be higher, as illustrated in Figure 2.
+This high similarity indicates that certain layers induce minimal changes to the hidden states, suggesting they contribute little to the model’s overall function.
+A detailed mathematical explanation for this phenomenon is provided in Appendix A.
+Which suggests that the deep layers of the model with pre-norm might not play a critical role in the overall function, and that the layers in large language models could be more redundant than expected, which motivates the layer-removal based pruning method we explore in the next section.
+2.2 Layer redundancy
+Table 1: Ablation of removing FFN and Attention of Llama2-7B-Base. We sample 100 instances from PG19 (Rae et al., 2019) to calculate PPL.
+Delete PPL
+None 7.60
+The whole last layer 13.37
+Attention of the last layer 7.65
+FFN of the last layer 12.35
+As discussed in the previous section, we speculate that the LLMs exhibit layer redundancy.
+To verify this, we assess the performance degradation caused by removing individual layers of two popular models, Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese.
+Figure 1 confirms our speculation, which reveals that some layers do not play a crucial role in LLMs, causing little degradation when omitting them individually.
+Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the last layer often being more critical.
+Notably, we found the last layer to be particularly important, aligning with findings from LLM Pruner (Ma et al., 2024).
+This observation contradicts our mathematical explanation in Appendix A which suggests that deeper layers tend to be more redundant.
+We posit that this discrepancy arises because the final FFN effectively functions as part of the token classifier and should be considered in conjunction with the language model head.To verify our hypothesis, we conducted further investigation, detailed in Table 1.
+The results show that within the last layer, the FFN component is crucial, while the Attention module is less significant.
+This finding supports our interpretation of the final layer’s importance.
 3 Methodology
-In this section, we present the methodological framework of our layer deletion approach for LLMs, elucidating the underlying principles and techniques employed.
-We first quantify the layer redundancy problem present in current prominent LLMs, such as LLaMA 2 and Baichuan 2.
-Then, we introduce a metric called Block Influence (BI), designed to assess the transformation of hidden states by each layer during LLM inference.
-Building upon BI, we apply layer deletion to LLMs, thereby reducing their inference cost without compromising their predictive accuracy or linguistic capabilities.
-2.1 Layer redundancy
-The predominant LLMs in current are primarily based on the Transformer (Vaswani et al., 2017).
-The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other.
-Transformer is a sequence to sequence mapping, which can be defined as $y=f(X,\theta)$, where $X\in R^{T\times n}$, $Y\in R^{T\times n}$, $T$ is the length of the sequence, $n$ is the vocabulary size, $\theta$ is the learnable parameters.
-The formal expression of an L-layer transformer is as follows:
-$\textbf{\footnotesize{X}}_{0}$ $=\textbf{\footnotesize{X}}\textbf{W}_{\text{emb}}$
-$\textbf{\footnotesize{A}}_{i+1}$ $=\text{ATTN}(\text{LN}(\textbf{\footnotesize{X}}_{i}))+\textbf{\footnotesize{X}}_{i}$
-$\textbf{\footnotesize{X}}_{i+1}$ $=\text{FFN}(\text{LN}(\textbf{\footnotesize{A}}_{i+1}))+\textbf{\footnotesize{A}}_{i+1}$
-Y $=\textbf{\footnotesize{X}}_{L}\textbf{W}_{\text{head}}$ (1)
-where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ (Chowdhery et al., 2023), $d$ is the hidden dim of the transformer.
-ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers.
-Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are.
-Previous works found that the Transformer possesses certain semantic capabilities in earlier layers (Hasan et al., 2021).
-In this work, we uncover a significant level of redundancy between the layers of the transformer.
-Layer Redundancy: A network with high redundancy should contain some redundant layers, which have a minimal impact on the network’s final performance.
-This may be due to these layers having homogenized functionalities compared to other layers in the network.
-We have found a high degree of redundancy in the current LLMs by omitting specific layers.
-(a) Llama2-7B-Base
-(b) Baichuan2-7B-Base
-Figure 3: Performance of removing single layer.
-We observed that in many cases, the network’s prediction of $Y$ remains largely unchanged.
-This phenomenon demonstrates a significant level of redundancy across the layers of the transformer.
-Figure 3 shows the perplexity and MMLU (Hendrycks et al., 2020) score of omitting a certain layer in the Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese.
-For more benchmark results, please refer to the section 3.
-Figure 3 reveals that some layers do not play a crucial role in LLMs.
-Dropping a certain layer from the LLMs may have a minimal effect on the final results, and the same holds for the perplexity.
-Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the final layer often being more critical.
-Figure 4: Illustration of our Block Influence(BI) metric. We omit LayerNorm for visual conciseness.
+In this section, we present the methodological framework of our layer removal approach for LLMs, elucidating the underlying principles and techniques employed.
+We begin by introducing Block Influence (BI), a novel metric designed to assess the hidden states transformation of each layer.
+Leveraging BI, we then detail our layer removal method.
 3.1 Layer importance
 As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers.
-Notably, the deeper layers demonstrate a higher level of redundancy.
-To remove these redundant layers, we need a model-intrinsic metric to measure the importance of the layer.
-An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons.
-In (Samragh et al., 2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers.
-To characterize the influence of a certain layer.
-We defined a new metric, Block Influence (BI), under the assumption that the more a transformer block changes the hidden states, the more influential this layer is.
-As Figure 4 shows, the BI score of $i^{th}$ blocks can be calculated as follows:
+To capture this, we introduce a new metric, Block Influence (BI), to measure the degree of transformation performed by each layer.
+The BI score of $i^{th}$ layer can be calculated as follows:
 $\text{BI}_{i}=1-\mathbb{E}_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{||X_{i,t}||_{2}||X_{i+1,t}||_{2}},$ (1)
-where $X_{i,t}$ means the $t^{th}$ row of $X_{i}$. Our empirical evidence supports that BI effectively reflects the importance of a layer. Figure 8 shows these different metrics in detail.
+where $X_{i,t}$ means the $t^{th}$ row of hidden states of $i^{th}$ layer.
+Lower BI score imply that $X_{i}$ and $X_{i+1}$ exhibit high cosine similarity, suggesting that the layer makes minimal transformations to the hidden states and is therefore less important.
+We plot the BI scores of a single layer and the PPL after removing it separately, as shown in the Figure 3.
+The results demonstrate a positive correlation between the BI score and the importance of a layer.
+(a) Llama2 7B
+(b) Baichuan2 7B
+Figure 3: The BI score of a layer and the PPL after removing the layer.
 3.2 Layer Removal
-The method we propose is straightforward: layer removal, which we simply referred as ShortGPT.
+Our goal is to obtain a pruned model that remains as close as possible to the original model.
+Since an LLM functions as a series of transformations applied to hidden states across its layers and we can determine the importance of each layer, we propose a straightforward pruning method: layer removal, which we refer to as ShortGPT.
 We delete certain layers in LLMs based on BI score.
-Figure 2 illustrates our pruning method.
 First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 (Rae et al., 2019).
@@
 Next, we calculate the BI score based on the collected hidden states.
-Finally, we sort layers in ascending order according to the BI, and delete the layers with the smaller importance.
+Finally, we sort layers in ascending order according to the BI, and delete the layers with the lower BI score.
 The number of layers to be deleted can vary to trade off the speed and performance.
-The details of our layer remove setting can be found in Appendix B
+The details of our layer removal setting can be found in Appendix D.
 4 Experiments
 4.1 Experimental Setup
-3.1.1 Models
+Models.
 To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B (Touvron et al., 2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B.
@@
 Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE (Su et al., 2024) positional embedding with ALiBi (Press et al., 2021).
-3.1.2 Benchmarks
-In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024) for evaluating large models.
-In addition, we also followed LaCo (Yang et al., 2024) to evaluate a wider dataset.
-MMLU (Hendrycks et al., 2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings.
-This makes benchmarks more challenging and similar to the way we evaluate humans.
-This benchmark covers 57 subjects including STEM, humanities, social sciences, etc.
-Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
-CMMLU (Li et al., 2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture.
-CMMLU covers 67 topics, from elementary school to university or professional level.
-Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
-CMNLI (Xu et al., 2020) is part of the Chinese language understanding assessment benchmark.
-It consists of two parts: XNLI and MNLI.
-HellaSwag (HeSw) (Zellers et al., 2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans.
-PIQA (Bisk et al., 2020) is a multi-choice question and answer dataset that focuses on daily scenarios.
-This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios.
-CHID (Zheng et al., 2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms.
-CoQA (Reddy et al., 2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers.
-BoolQ (Clark et al., 2019) is a question-answer dataset containing 15942 examples of yes/no questions.
-These problems occur naturally - they are generated in an environment that is silent and unconstrained.
-Race (Lai et al., 2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students.
-XSum(Hasan et al., 2021) is used to evaluate abstract single document summarization systems.
-The goal is to create a short, one-sentence new summary of what the article is about.
-C3 (Sun et al., 2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams.
-PG19 (Rae et al., 2019) is a long document dataset from books used to test the effectiveness of language modeling.
-3.1.3 Baselines
-To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo (Yang et al., 2024).
-For our method, we use PG19 for layer importance and perplexity calculation.
-LLMPru.
-(Ma et al., 2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
+Benchmarks.
+In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted comprehensive evaluation from five aspect: Reasoning: CMNLI (Li et al., 2024), HellaSwag (HeSw) (Zellers et al., 2019), PIQA (Bisk et al., 2020).
+Language: CHID (Zheng et al., 2019), WSC (Levesque et al., 2012).
+Knowledge: CommonSenseQA (CoQA) (Reddy et al., 2019), BoolQ (Clark et al., 2019).
+Examination: MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024).
+Understanding: Race-High/Middle (H/M) (Lai et al., 2017), XSum (Hasan et al., 2021), C3 (Sun et al., 2020) and PG19 (Rae et al., 2019).
+For more details, please refer to Appendix G
+Baselines.
+To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models, including:
+1) LLMPru (Ma et al., 2024), which adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
 LLMPru.
 applies post training to the pruned model, but for fair comparison, we do not apply post training to it.
-SliceGPT (Ashkboos et al., 2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
+2) SliceGPT (Ashkboos et al., 2024), which is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
 Specifically, they applied PCA to the hidden representation from shallow to deep layers, and incorporated the dimension reduction matrix into existing network parameters.
-LaCo (Yang et al., 2024) is a pruning method for large language models based on reducing layers.
+3) LaCo (Yang et al., 2024), which is a pruning method for large language models based on reducing layers.
 LaCo gradually merges similar layers from deep to shallow and sets a threshold to avoid continuously merging too many layers.
+For our evaluation, we use PG19 for layer importance and perplexity calculation. The models, baselines and evaluate benchmarks is the same as LaCo.
+Table 2: Comparison of pruning methods on multiple natural language benchmarks.
+The results of LLMPrun., SliceGPT and LaCo are reported from LaCo.
+The last column reports the relative performance retention.
 LLM Method Ratio Benchmarks Ave. Per.
 CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU
-Llama2-7B Dense 0.00% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92 44.52 100.00
-LLMPrun. 27.0% 34.33 56.46 71.22 25.25 36.54 42.51 55.20 22.56 22.35 11.51 25.64 23.33 25.25 32.84 73.76
-SliceGPT 26.4% 31.70 50.27 66.21 20.79 36.54 41.36 38.32 21.07 21.66 4.89 39.78 28.92 25.37 32.84 73.76
-LaCo 27.1% 34.43 55.69 69.80 36.14 40.38 45.70 64.07 22.61 23.61 15.64 39.67 26.45 25.24 38.41 86.28
-ShortGPT 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.71 32.25 35.17 0.67 39.62 43.96 32.25 42.60 95.69
-Llama2-13B Dense 0.00% 32.99 74.78 79.71 47.35 50.00 66.91 82.39 57.95 60.38 23.45 47.51 55.00 38.40 51.91 100.00
-LLMPrun. 24.4% 33.03 67.76 76.66 35.64 40.38 50.86 56.42 22.47 22.08 19.17 32.33 25.21 24.71 38.97 75.07
-SliceGPT 23.6% 29.82 55.71 69.04 19.31 36.54 47.26 37.86 23.41 24.03 5.27 41.92 37.14 25.79 34.84 67.11
-LaCo 24.6% 32.86 64.39 74.27 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 48.30 93.05
-ShortGPT 24.6% 33.00 66.64 73.45 36.61 50.00 58.64 62.48 58.35 60.17 17.59 46.90 54.69 38.38 50.53 97.34
-Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 26.96 24.09 20.82 64.55 53.87 56.95 53.63 100.00
-LLMPrun. 24.2% 32.28 53.66 71.82 69.80 53.85 47.83 61.19 21.96 22.28 15.98 41.64 24.93 25.69 41.76 77.87
-SliceGPT 22.2% 32.07 25.29 50.33 14.85 36.54 19.57 39.30 23.53 22.49 0.00 26.58 25.18 25.25 26.23 56.38
-LaCo 24.2% 33.00 52.28 68.50 76.24 42.31 47.26 56.15 28.99 27.72 12.03 50.85 31.53 31.24 42.93 80.05
-ShortGPT 24.2% 33.30 56.96 67.68 65.63 50.00 46.70 67.83 53.26 46.76 0.04 56.33 45.77 47.87 49.08 91.52
+Llama2-7B Dense 0.00% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92 47.78 100.00
+LLMPrun. 27.0% 34.33 56.46 71.22 25.25 36.54 42.51 55.20 22.56 22.35 11.51 25.64 23.33 25.25 34.78 72.79
+SliceGPT 26.4% 31.70 50.27 66.21 20.79 36.54 41.36 38.32 21.07 21.66 4.89 39.78 28.92 25.37 32.84 68.73
+LaCo 27.1% 34.43 55.69 69.80 36.14 40.38 45.70 64.07 22.61 23.61 15.64 39.67 26.45 25.24 38.41 80.39
+ShortGPT 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.71 32.25 35.17 0.67 39.62 43.96 32.25 41.24 86.31
+Llama2-13B Dense 0.00% 32.99 74.78 79.71 47.35 50.00 66.91 82.39 57.95 60.38 23.45 47.51 55.00 38.40 55.14 100.00
+LLMPrun. 24.4% 33.03 67.76 76.66 35.64 40.38 50.86 56.42 22.47 22.08 19.17 32.33 25.21 24.71 38.97 70.67
+SliceGPT 23.6% 29.82 55.71 69.04 19.31 36.54 47.26 37.86 23.41 24.03 5.27 41.92 37.14 25.79 34.85 63.20
+LaCo 24.6% 32.86 64.39 63.20 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 47.62 86.36
+ShortGPT 24.6% 33.00 66.64 73.45 36.61 50.00 58.64 62.48 58.35 60.17 17.59 46.90 54.69 38.38 50.53 91.64
+Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 52.63 51.04 20.82 64.55 53.87 56.95 57.67 100.00
+LLMPrun. 24.2% 32.28 53.66 71.82 69.80 53.85 47.83 61.19 21.96 22.28 15.98 41.64 24.93 25.69 41.76 72.41
+SliceGPT 22.2% 32.07 25.29 50.33 14.85 36.54 19.57 39.30 23.53 22.49 0.00 26.58 25.18 25.25 26.23 45.48
+LaCo 24.2% 33.00 52.28 68.50 76.24 42.31 47.26 56.15 28.99 27.72 12.03 50.85 31.53 31.24 42.93 74.44
+ShortGPT 24.2% 33.30 56.96 67.68 65.63 50.00 46.70 67.83 53.26 46.76 0.04 56.33 45.77 47.87 49.08 85.10
 Baichuan2-13B Dense 0.00% 33.21 71.10 78.07 86.51 50.00 65.6 77.89 67.27 68.94 25.02 65.64 59.50 61.30 62.31 100.00
 LLMPrun. 24.3% 33.80 53.57 71.82 72.77 37.50 38.82 56.54 21.17 21.61 13.67 39.89 23.19 25.18 39.20 62.91
-SliceGPT 22.8% 32.07 25.85 51.03 10.40 36.54 18.02 37.83 21.56 21.52 0.00 24.99 22.95 25.26 25.03 40.17
+SliceGPT 22.8% 32.07 25.85 51.03 10.40 36.54 18.02 37.83 21.56 21.52 0.00 24.99 22.95 25.26 25.23 40.49
 LaCo 24.7% 33.03 60.71 68.88 76.73 44.23 55.45 62.35 56.92 57.80 12.32 61.10 51.35 53.65 53.43 85.75
-ShortGPT 24.7% 32.81 60.55 71.60 80.17 47.13 54.30 62.54 55.77 56.41 15.14 60.16 52.11 58.86 54.43 87.33
-Table 1: Comparison of pruning methods on multiple natural language benchmarks. The results of LLMPrun., SliceGPT and LaCo are reported from LaCo.
+ShortGPT 24.7% 32.81 60.55 71.60 80.17 47.13 54.30 62.54 55.77 56.41 15.14 60.16 52.11 58.86 54.43 87.35
 4.2 Main Results
-To validate the efficacy of our proposed method, we conducted comparative experiments against benchmark and baseline techniques commonly employed in large language model evaluation.
+To validate the efficacy of our proposed method, we conducted comparative experiments against baseline techniques commonly employed in large language model evaluation.
 Considering the current structured pruning methods generally reduce parameters by no more than 30%, we performed experiments with approximately 1/4 of the parameters pruned.
@@
 Additional experiments exploring different parameter reduction proportions will be discussed in the subsequent section.
-The results demonstrate that the performance of the model pruned by our method significantly surpasses that of the baseline method, maintaining most of the large language model’s capabilities, such as reasoning, language understanding, knowledge retention, and examination performance, to the greatest extent possible.
-Furthermore, we note that the approach of reducing the number of layers (ShortGPT/LaCo) outperforms the method of reducing the embedding dimensions (LLMPru./SliceGPT), implying that the model exhibits more redundancy in-depth than in width.
+The results demonstrate that the performance of the model pruned by our method significantly surpasses that of the baseline methods, maintaining most of the large language model’s capabilities.
+Furthermore, we note that the approach of reducing the number of layers (ShortGPT/LaCo) outperforms the method of reducing the embedding dimensions (LLMPru./SliceGPT), implying that the model exhibits more redundancy in depth than in width.
 Further experimental analysis will be presented in the ensuing section.
-4 Analysis
-4.1 Varying prune ratio
-Figure 5: Pruning by different metrics, with increasing pruning ratio.
-As a pruning method, we further validated the effects of different pruning ratios on model performance.
-Experiments were conducted on the Llama2-7B-Base and Baichuan2-7B-Base models, observing the Perplexity and MMLU.
-The pruning ratios ranged from 0% to 97%, employing the strategy of deleting layers based on the importance proposed in this paper, resulting in models with 1 to 32 layers.
-The results in Figure 5 indicate that as the pruning ratio increases, the performance of the model declines.
-However, the MMLU score drops significantly at a certain layer, which may imply the existence of certain special layers within the network that play a critically important role.
-4.2 Depth Redundancy v.s. Width Redundancy
-Figure 6: Redundancy on width, removing a single head from model. The x-axis is the head id.
-In the preceding sections, we analyzed and explored the redundancy of large language models (LLMs) in terms of depth (layers).
-However, we also investigated redundancy in terms of width, specifically focusing on the attention heads.
-Our method involves sequentially removing each head from the multi-head attention mechanism and adjusting the parameters of the attention output projection to ensure the output shape of the Attention block remains unchanged.
-The Feed-Forward Network (FFN) remains unaltered.
-Similar to the approach in Section 2.1, we observe the impact of each head’s removal on the final performance of the network.
-Figure 6 illustrates the changes in Perplexity and MMLU scores after each head is removed.
-The results indicate that LLMs exhibit a high degree of redundancy in width, comparable to their redundancy in depth.
-However, this width redundancy does not follow any discernible pattern and varies across different models.
-We hypothesize that this variation may be attributed to the symmetry among heads.
-4.3 Redundancy on non-transformer LLM
-Figure 7: Redundancy on RWKV, removing single layers on RWKV-7B.
-We further investigate whether the redundancy observed in depth is a result of the Transformer’s architecture.
-Given that the vast majority of popular LLMs are based on the Transformer structure.
-We selected the RWKV-7B model proposed in (Peng et al., 2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent.
-We analyzed this model using the methodology from 2.1.
-Figure 7 shows the redundancy of the RWKV-7B 111 We use rwkv-v5-world-7B from https://huggingface.co/BlinkDL/rwkv-5-world model.
-Through this figure, we can observe that the RWKV-7B model also exhibits a high level of redundancy.
-This may suggest that the redundancy is universal across LLMs.
-4.4 Importance metric
-Figure 8: Different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer
-The overarching concept of our method is to rank layers by importance and remove the less significant ones.
-The definition of importance metrics has a profound impact on the outcome.
-In this section, we define and compare various important metrics:
-Sequential: The importance is directly proportional to the sequence order, with shallower layers being more important.
-This can be implemented by taking the negative of each layer’s index as the importance metric.
-Reverse-order/Norm: This metric posits that importance is inversely proportional to the sequence order.
-It assigns higher importance scores to the shallower layers Inspired by Figure 3.
-This method gives the same order as measuring importance by hidden states norm mentioned in LABEL:layerimportacne.
-Here, the importance metric is the layer’s index itself.
-Relative Magnitude: Proposed in (Samragh et al., 2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
-BI: BI mentioned in previous section 2.2.
+In Table 2, we fully adopted the benchmark, model, and pruning ratio in the LaCo paper.
+In order to make a more fair comparison with LLMprun.
+and SliceGPT, we compared them with the same benchmark, model, and pruning ratio in their original paper.
+The experimental results are shown in Appendix C.
+Consistent with our findings in Table 2, these experiments further demonstrate the significant layer redundancy present in existing large language models, and ShortGPT achieves superior performance compared to other pruning methods.
+The results show that coarse-grained pruning methods, such as removing entire layers, often outperform fine-grained approaches like Slice GPT or LLM Pruner.
+We speculate that the reason is that the large language model is actually very robust, as shown in Figure 1, removing any deep layer individually actually has very little impact on the final output, which means it is difficult to define the importance of a finer grained module and perform pruning.
+4.3 Varying metric and pruning ratio
+Figure 4: Comparison of different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer.
+The core principle of our method is to rank layers by their importance and remove the less significant ones.
+The choice of importance metric significantly influences the outcome.
+In this section, we define and compare several different importance metrics:
+  • Sequential: The importance is directly proportional to the sequence order, with shallower layers being less important.
+This can be implemented by assigning the negative value of each layer’s index as its importance metric.
+  • Norm/Reverse-order: This metric posits that importance is inversely proportional to the sequence order.
+It assigns higher importance scores to the shallower layers.
+This method gives the same order as measuring importance by hidden states norm as Figure 4 shows.
+  • Relative Magnitude: Proposed in Samragh et al. (2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
+  • BI: we calculate the BI score mentioned in Section 3.1 as importance metric.
 Figure 4 demonstrates the different metrics.
@@
 Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics.
-The method of (Samragh et al., 2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent.
-It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation.
-This also indicates the limitations of the current LLMs evaluation methods.
+The method of Relative Magnitude is highly competitive, indicating that relative values can also reflect the importance to some extent.
+It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor.
+Figure 5: Performance of MMLU and perplexity when we prune by different metrics, with increasing pruning ratio. We can see that as the pruning ratio increases, the performance of the model declines.
+As a pruning method, we further validated the effects of different pruning ratios on model performance.
+Experiments were conducted on the Llama2 and Baichuan2 models, observing the Perplexity and MMLU.
+The results for Llama2, as shown in Figure 5, demonstrate that the model’s performance generally declines as the pruning ratio increases.
+However, we observe a notable phenomenon: the MMLU score exhibits a sharp drop at a specific layer.
+This sudden decrease suggests the presence of certain critical layers within the network that play a particularly important role in maintaining performance.
+Similar patterns are observed in the Baichuan2 model, as illustrated in Appendix B.
+4.4 Redundancy on non-transformer LLM
+To determine whether the observed depth redundancy is specific to the Transformer architecture, we extended our investigation to include two popular non-Transformer models, RWKV-7B111 We use rwkv-v5-world-7B from https://huggingface.co/RWKV/v5-Eagle-7B-HF (Peng et al., 2023) and Mamba-2.8B  [^2] We take the model from https://huggingface.co/state-spaces/mamba-2.8b-hf (Gu & Dao, 2023).
+Our experiments revealed that these models also exhibit resilience to layer removal, maintaining performance despite the elimination of certain layers.
+This finding suggests that the redundancy phenomenon may not be unique to Transformer-based models, but rather a common characteristic across current large language models.
+Table 3 shows that our method is applicable and effective for both Mamba and RWKV models, suggesting that the redundancy is universal across current LLMs.
+However, it is worth noting that the RWKV model appears less redundant than Mamba and Transformer models, which warrants further investigation.
+Table 3: ShortGPT pruning on RWKV and Mamba.
+Model Pruning ratio CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU Ave. Per.
+Mamba-2.8B 0% 35.97 61.84 75.52 35.56 49.69 56.35 60.67 24.9 25.3 15.03 42.08 26.29 25.32 41.12 100.00
+10.9% 32.95 59.71 73.01 32.52 49.28 52.66 51.41 24.27 25.21 14.95 41.1 26.01 25.00 39.08 95.04
+20.3% 31.29 55.69 69.64 29.12 48.36 48.32 62.2 23.61 23.61 14.71 41.59 25.69 25.37 38.36 93.29
+25% 29.96 52.38 68.77 26.02 48.26 44.96 62.2 23.67 23.26 14.00 40.71 24.32 24.89 37.18 90.42
+31.3% 28.25 47.02 64.91 21.38 49.69 44.96 62.17 21.87 22.77 13.77 40.44 24.48 24.77 35.59 86.55
+RWKV-7B 0% 32.07 65.98 77.09 85.36 50.00 62.65 62.72 38.56 45.47 16.5 57.97 31.85 28.54 50.37 100.00
+9.4% 32.6 56.41 73.94 78.12 50.00 49.55 62.35 25.9 25.77 9.57 54.68 27.29 25.03 43.94 87.23
+18.8% 32.11 49.47 71.55 65.63 50.00 40.54 61.19 22.04 23.75 8.13 49.15 26.35 25 40.38 80.17
+25% 32.41 39.73 65.13 52.6 50.00 29.65 60.92 22.56 21.59 12.02 41.86 25.52 25.08 36.85 73.16
+28.1% 33.11 32.22 60.01 32.47 50.1 28.34 60.85 22.27 21.31 10.43 37.81 25.64 25.15 33.82 67.14
 4.5 Orthogonal to Quantization
 In this section, we show that our method is orthogonal to quantization methods.
-We apply our method to a model  [^2] We take the model from https://huggingface.co/TheBloke/Llama-2-7B-GPTQ quantized by GPTQ algorithm.
-Table 2 shows that our method is compatible with the quantization-like method, which could reduce the footprint further.
-Model Ratio/Layer Perplexity MMLU
-Baseline 0%/32 8.4999 37.99
-ShortGPT 27.1%/23 42.6951 36.69
+We apply our method to Llama2-7B  [^3] We take the model from https://huggingface.co/TheBloke/Llama-2-7B-GPTQ quantized by GPTQ algorithm.
+Table 4 shows that our method is compatible with the quantization-like method.
+In addition, we compared the performance of applying pruning before quantization  [^4] We use GPTQ algorithm for quantization from https://github.com/AutoGPTQ/AutoGPTQ.
+The results shown in the Table 5 further indicates that quantization and ShortGPT are orthogonal operations.
 Table 4: Layer removal results on Llama2-7B-Base-GPTQ.
+Model Ratio/Layer Perplexity MMLU Throughput (speed up)
+Baseline 0%/32 8.03 43.17 4331.23 Token/s (1.00x)
+3.1%/31 8.37 42.88 4399.31 Token/s (1.02x)
+9.4%/29 9.44 42.31 4602.26 Token/s (1.06x)
+ShortGPT 12.5%/28 10.24 41.62 4680.68 Token/s (1.08x)
+15.6%/27 11.42 43.17 4756.94 Token/s (1.10x)
+25.0%/24 22.29 41.68 5045.59 Token/s (1.16x)
+27.1%/23 40.78 43.35 5146.99 Token/s (1.19x)
+Table 5: Performance comparison of different methods
+Method MMLU CMMLU
+Llama2-7B-Baseline 45.4 32.9
+4-bit quantization 44.9 32.5
+Layer removal (27.1%) 44.0 32.3
+4-bit quantization then layer removal 42.4 31.0
+Layer removal then 4-bit quantization 41.2 30.5
+4.6 Post training to restore performance
+To mitigate the performance loss resulting from layer removal, we explored post-training strategies inspired by Chen et al.
+(2024).
+Our approach comprised two key steps: 1)Replacement: We substituted the removed layers with lightweight Multi-Layer Perceptron (MLP) modules.
+2)Retraining: We subsequently retrained the modified model.
+The results in Table 6 demonstrate the potential of post-train in recover performance loss.
+Appendix F list the training details.
+Table 6: Post-train Llama2-7B to restore performance.
+Method Avg. Ratio CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU
+Dense 47.78 0% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92
+ShortGPT 41.22 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.41 32.25 35.17 0.67 39.62 43.96 32.25
+ShortGPT+post-train 43.16 24.0% 32.99 54.83 68.12 31.82 51.37 58.32 72.36 34.18 34.68 4.89 40.37 44.47 32.73
 5 Limitation
-Although our method demonstrates strong competitiveness compared to current pruning methods, with some benchmarks even suggesting that deleting certain layers has no impact on the final result, this does not imply that layer removal comes without drawbacks.
-Due to the limitations of benchmarking, current evaluations may not fully capture the impact of layer removal on model performance.
-Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks.
-On benchmarks such as GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero.
-Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal.
-Post-training techniques could potentially further mitigate performance losses, which is an area of ongoing research.
+Although our method demonstrates strong competitiveness compared to current pruning methods, there are some phenomena that have not been explained.
+Our experiments reveal that the negative effect of layer removal is more significant on generative tasks compared to multiple-choice tasks.
+When we remove 25% layers from Llama2-7B or Baichuan2-7B, the performance in generative tasks such as XSum and C3 deceases to nearly zero, although the performance decline was not as significant on the larger model of the 13B.
+We speculate that compared to multiple-choice tasks, generative tasks face the problem of accumulated errors and large model is more robust than small one.
+The reasons behind it still need to be explored.
+The post-training techniques discussed in Section 4.6 have the potential to mitigate this issue and warrant further exploration.
 6 Related works
-To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into four categories: model pruning, knowledge distillation, quantization, and low-rank factorization.
+To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into two categories: model pruning and quantization.
+Besides, there are some works aim to study the redundancy of model which is essential for compressing models.
 Model pruning: model pruning (LeCun et al., 1989; Han et al., 2015) is a classic and effective method of reducing model redundancy modules to compress models.
 The model pruning methods mainly include unstructured pruning and structured pruning.
-The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar and Alistarh, 2023) and LoRAPrune (Zhang et al., 2023).
+The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar & Alistarh, 2023) and LoRAPrune (Zhang et al., 2023).
 However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition.
@@
 Compared to the previous method, our method is a simple and efficient structured pruning method.
-Knowledge distillation: Knowledge distillation (KD) (Hinton et al., 2015; Gou et al., 2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network.
-Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results (Li et al., 2022).
-In addition, when we can access transparent teacher models, the result can be improved more.
-MiniLLM (Gu et al., 2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution.
-DistlLLM (Ko et al., 2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs.
-However, compared to model pruning, this method often requires higher computational resources.
-Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models.
+Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022; 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models.
 Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms.
@@
 Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
-Low-rank factorization: low-rank decomposition (Cheng et al., 2017; Povey et al., 2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions.
-TensorGPT (Xu et al., 2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices.
-Recently, SliceGPT (Ashkboos et al., 2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
-Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole and Morgan, 1997).
+Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole & Morgan, 1997).
 In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy.
@@
 7 Conclusion
-This work has introduced a novel approach to pruning Large Language Models (LLMs) based on layer redundancy and an "importance" metric defined as attention entropy.
-Our study reveals a significant degree of layer-wise redundancy in LLMs, indicating that certain layers contribute minimally to the overall network function and can thus be removed without substantially compromising model performance.
-By employing a straightforward layer removal strategy guided by the calculated importance of each layer, we have demonstrated that it is possible to maintain up to 95% of an LLM’s performance while reducing the model’s parameter count and computational requirements by approximately 25%.
-This achievement not only surpasses previous state-of-the-art pruning methods but also underscores the potential for further optimizations in LLM deployment strategies.
-Our findings suggest that the redundancy inherent in LLMs is largely depth-based rather than width-based, highlighting an avenue for future research into the structural efficiency of neural networks.
-Moreover, our pruning approach, characterized by its simplicity and effectiveness, is compatible with other compression techniques such as quantization, offering a composite pathway to model size reduction that is both significant and versatile.
-The implications of our research extend beyond academic interest, offering practical solutions for deploying advanced LLMs in resource-constrained environments.
-By enabling more efficient model architectures without the need for extensive retraining, our pruning method facilitates broader access to cutting-edge AI capabilities across a variety of platforms and devices.
-In conclusion, our investigation into layer redundancy and the development of an importance-based pruning strategy represent a meaningful advancement in the optimization of Large Language Models.
-As the demand for sophisticated AI tools continues to grow, approaches like ours will play a crucial role in making these technologies more accessible and sustainable.
-Future work will focus on refining our understanding of model redundancy and exploring additional methods for enhancing the efficiency of neural network models.
-Appendix A Detailed Strategy for Layer Removal
-Table 3: Strategies for Layer Removal in Models.
-Strategy
-Description
-Sequential
-Layers are removed sequentially from the beginning of the model. The process starts with layer 0 and progressively includes more layers for removal (e.g., {0}, {0, 1}, …).
-Reverse-order
-This strategy involves starting from the model’s final layer and progressively removing layers in reverse order (e.g., {-1}, {-1, -2}, …).
-Relative Magnitude
-Layers are removed in ascending order based on their Relative Magnitude values.
-The removal process accumulates layers from those with the smallest to the largest values, mirroring the sequential strategy’s accumulation method.
-BI (Block Influence)
-Follows a similar accumulation approach as the Sequential strategy, but layers are ordered and removed according to their BI values, starting from the lowest and moving to the highest.
-Appendix B Benchmark Setup for Removed Layers
-Before delineating the specific layer removal configurations in the benchmarked models, a concise overview of the structural modifications is warranted.
-The Llama-2-7B and Baichuan-2-7B models, initially designed with 32 layers, have been streamlined by excising 9 layers, corresponding to a 28% reduction of the total layer framework.
-The 13B variants, originally constituted of 40 layers, experienced a decrement of 10 layers, marking a 25% diminution.
-The selection criterion for layer removal was based on ascending Block Influence (BI) values, ensuring a data-driven approach to minimization.
-This premise underpins the benchmark configurations.
-Table 4: Setup of Removed Layers for Benchmark Models
+In this work, we uncovered the significant layer-wise redundancy of LLMs, Our research demonstrates that certain layers contribute minimally to overall network functionality and can be removed without substantially compromising model performance.
+Based on our observation, We introduce Block influence to quantify the importance of each layer and propose a simple and straightforward pruning method: layer removal.
+Our experiments demonstrates that it is possible to maintain up to approximately 90% of a LLM’s performance while reducing the model’s parameter amount and computational requirements by approximately 25%.
+Besides, our method is orthogonal to quantization methods and can be further improved by continual training.
+We hope that our work can provide some insight for future model compression techniques.
+Moreover, our work suggests potential avenues for improving the efficiency of model training by reducing inherent redundancy in the future.
+Appendix A Mathematical explanation for why pre-norm brings high similarity
+We provide a simple explanation here about how pre-norm leads to high deep similarity in this section, here we adopt RMSNorm (Zhang & Sennrich, 2019) for convenient, which is also the popular pre-norm used in many recent LLMs, such as Llama and Mamba.
+Lemma 1
+(Xiong et al., 2020) At initialization, for the Pre-LN Transformer, $(1+\frac{L}{2})d\leq\mathbb{E}(||x_{L,i}||^{2}_{2})\leq(1+\frac{3L}{2})d$ for all $L>0$ and $i$.
+Expectations are taken over the input and the randomness of initialization, where the hidden state of $L^{th}$ layer is $x_{L}$.
+From Lemma 1, the hidden state of the pre-norm model will continuously increase as the number of layers increases.
+And under the assumption of each component of $x_{l}$ has a mean of 00, we can obtain $||x_{L}||=\Theta(\sqrt{L})$.
+Then we consider $x_{L+1}=x_{L}+f_{L}(x_{L},\theta_{L})$, where $f_{L}$ is a operation such as Attention or MLP, $\theta_{L}$ is learnable parameters.
+Then $f_{L}(x_{L},\theta_{L})=O(1)$ respect to $L$, for Attention as example, $||f_{L}(x_{L},\theta_{L})||=||(softmax(Q^{T}K)X_{L}/||X_{L}||\cdot(\sigma_{rms}))W_{v}W_{q}||=O(||\sigma_{rms}||||W_{v}||||W_{o}||)=O(1)$ respect to $L$.
+Then we can get:
+cos $\textit{cos similarity}(X_{L+1},X_{L})$ $=\frac{x_{L+1}x_{L}}{||x_{L+1}||||x_{L}||}=\frac{||x_{L}||^{2}}{||x_{L+1}||||x_{L}||}+\frac{f_{L}(x_{L},\theta)x_{L}}{||x_{L+1}||||x_{L}||}$ (2)
+$\geq\frac{||x_{L}||^{2}}{||x_{L+1}||||x_{L}||}-\frac{||f_{L}(x_{L},\theta)||||x_{L}||}{||x_{L+1}||||x_{L}||}$ (3)
+$=\frac{||x_{L}||}{||x_{L+1}||}-\frac{||f_{L}(x_{L},\theta)||}{x_{L+1}}=\Theta(\sqrt{\frac{L}{L+1})}-O(\sqrt{\frac{1}{L+1}})$ (4)
+This means that as the number of layers $L$ increases, the similarity between the input and output of the layer will be high.
+This means that the role of $f_{L}$ may be relatively small, and removing it from the network may have a relatively small impact to the model.
+Although the above theoretical analysis is only for randomly initialized models, this phenomenon that deep layer has similar input and output exists in both our own trained models shown in Figure 2 and existing models in Figure 4.
+Appendix B Layer Removal on Baichuan2-series Model
+Figure 6: Pruning by different metrics on Baichuan2-series model.
+Appendix C A Fair comparison with SliceGPT and LLMprun.
+In Table 2, we fully adopted the benchmark, model, and pruning ratio in the LaCo’s paper.
+For a fair comparison with LLM pruner and SliceGPT, we do the same experiments in the original paper of LLM pruner and SliceGPT.
+The results is provided in Table 7 and Table 8.
+We take the same benchmarks, models and pruning ratio as the corresponding original paper.
+The results demonstrate that our method is highly competitive.
+[b]
+Table 7: Comparison between ShortGPT and LLM-pruner. The Table is corresponding to the Table 1 of LLM pruner(Zhang et al., 2023).
+Model Pruning ratio Method BoolQ PIQA Hellaswag Winogrande Arc-e Arc-c OBQA Avg.
+Llama-7B Ratio=0% Baseline 73.18 78.35 72.99 67.01 67.45 41.38 42.4 63.25
+Ratio=20% LLM-pruner 59.39 75.57 65.34 61.33 59.18 37.12 39.80 56.82
+Ratio=21.9 % ShortGPT 68.26 72.28 61.7 63.77 60.22 39 41.6 58.12
+Llama-13B Ratio=0% Baseline 68.47 78.89 76.24 70.09 74.58 44.54 42.00 64.97
+Ratio=20% LLM-pruner 67.68 77.15 73.41 65.11 68.35 38.4 42.4 61.79
+Ratio=20% ShortGPT 68.41 76.36 72.9 67.4 68.62 39.2 41 61.98
+[b] Model Pruning ratio Method PIQA Hellaswag Winogrande Arc-e Arc-c Avg.
+Llama-2-7B 0% Baseline 79.11 75.99 69.06 74.58 46.25 69 20% SliceGPT 71.87 58.1 63.04 69.87 43.09 63.45 25% SliceGPT 68.55 58.1 62.04 57.46 35.07 56.15 30% SliceGPT 66.1 52.69 56.82 35.07 56.82 56.15 21.9% ShortGPT 72.76 66.39 66.27 59.39 39.85 60.93 25% ShortGPT 70.53 62.68 64.7 58.39 39.51 59.16 31.6% ShortGPT 67.87 62.19 64.38 56.57 40.86 58.37 Llama-2-13B 0% Baseline 80.47 79.39 72.22 77.48 49.23 71.76 20% SliceGPT 71.87 69.38 63.04 69.87 43.09 63.45 25% SliceGPT 68.55 67.48 58.1 62.5 37.88 58.9 30% SliceGPT 66.1 65.11 52.69 56.82 35.07 55.16 20% ShortGPT 76.95 74.67 71.14 69.56 45.63 67.59 25% ShortGPT 74.39 71.65 70.98 67.09 43.93 65.61 30% ShortGPT 72.11 71.93 67.19 61.09 40.88 62.64 Llama-2-70B 0% Baseline 82.7 83.84 77.98 80.98 57.34 76.57 20% SliceGPT 76.61 72.98 74.92 80.51 55.2 72.34 25% SliceGPT 74.92 68.74 74.92 77.9 51.71 69.75 30% SliceGPT 72.31 63.69 73.4 51.71 47.61 66.11 20% ShortGPT 76.02 78.87 71.69 76.02 52.95 71.68 25% ShortGPT 73.2 76.72 71.85 73.2 49.9 69.79 30% ShortGPT 74.44 75.31 72.33 74.44 49.22 69.4
+Table 8: Comparison between ShortGPT and SliceGPT. The Table is corresponding to the Table 7 of SliceGPT(Ashkboos et al., 2024).
+Appendix D Detailed Strategies for Layer Removal
+We list the details of different layer removal strategies in Table 10. The concrete removed layers by ShortGPT in Table 2 are listed in Table 9
+Table 9: Setup of Removed Layers for Benchmark Models.
 Model Removed Layers
@@
 Baichuan-2-13B 32, 31, 33, 30, 34, 29, 28, 35, 27, 26
-Appendix C Layer Removal on Baichuan2-series Model
-Figure 9: Pruning by different metrics on Baichuan2-series model.
-Appendix D Layer Removal on HumanEval.
-We also found that layer removal has a significant impact on the performance of the model in generative tasks.
-Specifically, we conducted experiments on the Baichuan2-13B model on HumanEval (Chen et al., 2021), which is a dataset used to evaluate the performance of code generation models.
-The results are shown in Figure 10.
-After removing 10 layers, the performance on HumanEval almost completely degraded.
-We speculate that it may be due to significant cumulative errors in generative tasks.
-Figure 10: Performance degragation on generation task.
+Table 10: Strategies for Layer Removal in Models.
+Strategy Description
+Sequential Layers are removed sequentially from the beginning of the model. The process starts with layer 0 and progressively includes more layers for removal (e.g., {0}, {0, 1}, …).
+Reverse-order This strategy involves starting from the model’s final layer and progressively removing layers in reverse order (e.g., {-1}, {-1, -2}, …).
+Relative Magnitude Layers are removed in ascending order based on their Relative Magnitude values.
+The removal process accumulates layers from those with the smallest to the largest values, mirroring the sequential strategy’s accumulation method.
+BI (Block Influence) Follows a similar accumulation approach as the Sequential strategy, but layers are ordered and removed according to their BI values, starting from the lowest and moving to the highest.
+Appendix E Setup for training post-norm model and pre-norm model
+We have listed the specific training settings for pre norm and post norm in Table 11.
+Table 11: Training Parameters.
+Parameter Value
+Global Batch Size 2048
+Sequence length 4096
+Precision bf16
+Learning Rate Scheduler cosine
+Max Learning Rate 4e-4
+Min Learning Rate 5e-5
+Warm-up steps 3000
+Training Tokens 200B
+Weight Decay 0.1
+Adam Beta1 0.9
+Adam Beta2 0.98
+Gradient Clip 1.0
+Tokenizer Llama2
+Layers 32
+Hidden state 2048
+Attention heads 32
+Head dim 64
+FFN size 5504
+Activation function Silu
+Appendix F post-training settings
+We replace the removed layer with a lightweight gated MLP layer with hidden size = 2048. Table 12 show the post training settings.
+Table 12: Post-training Parameters.
+Parameter Value
+Global Batch Size 2048
+Sequence length 4096
+Precision bf16
+Learning Rate Scheduler cosine
+Max Learning Rate 2e-5
+Min Learning Rate 1e-5
+Warm-up steps 3000
+Training Tokens 50B
+Weight Decay 0.1
+Adam Beta1 0.9
+Adam Beta2 0.98
+Gradient Clip 1.0
+Appendix G Evaluation Benchmarks
+In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU Hendrycks et al.
+(2020), CMMLU Li et al.
+(2024) for evaluating large models.
+In addition, we also followed LaCo Yang et al.
+(2024) to evaluate a wider dataset.
+MMLU Hendrycks et al.
+(2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings.
+This makes benchmarks more challenging and similar to the way we evaluate humans.
+This benchmark covers 57 subjects including STEM, humanities, social sciences, etc.
+Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
+CMMLU Li et al.
+(2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture.
+CMMLU covers 67 topics, from elementary school to university or professional level.
+Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
+CMNLI Xu et al.
+(2020) is part of the Chinese language understanding assessment benchmark.
+It consists of two parts: XNLI and MNLI.
+HellaSwag (HeSw) Zellers et al.
+(2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans.
+PIQA Bisk et al.
+(2020) is a multi-choice question and answer dataset that focuses on daily scenarios.
+This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios.
+CHID Zheng et al.
+(2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms.
+CoQA Reddy et al.
+(2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers.
+BoolQ Clark et al.
+(2019) is a question-answer dataset containing 15942 examples of yes/no questions.
+These problems occur naturally - they are generated in an environment that is silent and unconstrained.
+Race Lai et al.
+(2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students.
+XSumHasan et al.
+(2021) is used to evaluate abstract single document summarization systems.
+The goal is to create a short, one-sentence new summary of what the article is about.
+C3 Sun et al.
+(2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams.
+PG19 Rae et al.
+(2019) is a long document dataset from books used to test the effectiveness of language modeling.
+Appendix H Hardware Environment
+The platform we use to experiment is GPU heterogeneous platform. The hardware of our platform is shown in Table 13
+Table 13: Setup of Removed Layers for Benchmark Models.
+Name Details
+CPU 2x Intel(R) Xeon(R) Gold 6430 CPU @ 2.1GHz
+GPU 8x NVIDIA A100-80GB Tensor Core GPU